2. A bag contains 13 marbles: 6 red ones, 4 blue ones, and 3 yellow ones. Suppose that two marbles are selected without replacement.
(a) What is the probability that a red marble, and then a yellow marble (in that order) are selected?
(b) What is the probability that neither of the two marbles is blue?

Answer by stanbon(75887)

A bag contains 13 marbles: 6 red ones, 4 blue ones, and 3 yellow ones.
Suppose that two marbles are selected without replacement.
--------------
(a) What is the probability that a red marble, and then a yellow marble (in that order) are selected?
There are 13*12 = 156 ordered pairs to select.
6*4 = 24 of them are red then blue
----
P(red then blue) = 24/156
======================================
(b) What is the probability that neither of the two marbles is blue?
There are 156 pairs
9*8 = 72 of them contain no blue
----
P(no blue) = 72/156
======================
Cheers,
Stan H.
=================